OpenStudy (anonymous):

The distance between two cities on a map is measured to be 4 inches. The actual distance between the cities is 448 miles. What is the scale of the map?

OpenStudy (nurali):

1/4 = x/448 448(1/4) = 448(x/448) 112 = x 1in = 112 miles

OpenStudy (anonymous):

thanks....

OpenStudy (nurali):

My Pleasure.